XYZ Construction Company is investigating the purchase of a new dump truck. Interest is 9%. The cash flows for two likely models are as follows:

Model First cost annual operating cost annual income salvage value life
A $50000 $2000 $9000 $10000 10 yr
B $80000 $1000 $12000 $30000 10 yr

a) Using PW analysis, decide which truck the firm should buy, and explain why.

b) Before the construction company can close the deal, the dealer sells out of Model B and can not get any more. What should the firm do now, and why?
Answer #1

Present worth Model A

PWA = -50000 -2000(P/A, 9%, 10) +9000(P/A, 9%,10) + 10000(P/F, 9%, 10)

= -50000 -2000× 6.418 + 9000×6.418 + 10000×0.4224

= -50000 - 12836 + 57762 + 4224

= - $ 850

Present worth of B

PWB = -80000 -1000(P/A, 9%,10) + 12000(P/A, 9%, 10) +30000(P/F, 9%, 10)

= -80000 -1000×6.418 + 12000×6.418 +30000×0.4224

= -80000 - 6418 + 77016 + 12672

= $3270

Present worth of Model B is greater than the present worth of A. Hence model B must be selected.

b. Present worth of A is negative hence it must be rejected.

Since PW of A is negative thus NPV will be negative that is the company is not able to earn what it invested.